# Stochastic Momentum Methods for Non-smooth Non-Convex Finite-Sum Coupled Compositional Optimization

### v1

@@ Appendix A Proofs of Technical Lemmas
Under Assumption 3.2(A1), $F(\cdot)$ in (1) is $\rho_{F}$-weakly convex with $\rho_{F}=L_{g}C_{f}+\rho_{f}C_{g}^{2}$. If $g_{i}$ is $L_{g}$ smooth, then $F_{\lambda}(\mathbf{w})$ is $L_{F}$-smooth with $L_{F}=\frac{C_{g}^{2}(2-\lambda\rho_{f})}{\lambda-\lambda^{2}\rho_{f}}+C_{f}L_{g}$.
@@
Consider any $\mathbf{w}$ and $\mathbf{w}^{\prime}$ in $\mathbb{R}^{d}$, any $i\in\{1,\dots,n\}$ and any $\mathbf{v}_{i}\in\partial f_{i}(\mathbf{t}_{i})$ at $\mathbf{t}_{i}=g_{i}(\mathbf{w})$. By Assumption 3.2(A1), $\mathbf{v}_{i}^{\top}g_{i}(\mathbf{w})$ is $\|\mathbf{v}_{i}\|L_{g}$-smooth in $\mathbf{w}$ and $f_{i}$ is $\rho_{f}$-weakly convex, we have
@@
$\geq f_{i}(g_{i}(\mathbf{w}))+\mathbf{v}_{i}^{\top}\nabla g_{i}(\mathbf{w}^{\prime})(\mathbf{w}^{\prime}-\mathbf{w})-\frac{\|\mathbf{v}_{i}\|L_{g}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2}-\frac{\rho_{f}C_{g}^{2}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2}$
$\geq f_{i}(g_{i}(\mathbf{w}))+\mathbf{v}_{i}^{\top}\nabla g_{i}(\mathbf{w}^{\prime})(\mathbf{w}^{\prime}-\mathbf{w})-\frac{C_{f}L_{g}+\rho_{f}C_{g}^{2}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2},$
where the second and the last inequalities hold because of lipschitzness of $f_{i}$ and $g_{i}$ from Assumption 3.2, the third because $\mathbf{v}_{i}^{\top}g_{i}(\mathbf{w})$ is $\|\mathbf{v}_{i}\|L_{g}$-smooth in $\mathbf{w}$.
By the $\rho_{f}$-weak convexity of $f_{i}$ and Proposition EC.1. in [36], we know that $f_{i,\lambda}$ is $L_{f_{\lambda}}:=\frac{2-\lambda\rho_{f}}{\lambda-\lambda^{2}\rho_{f}}$-smooth. Since $\nabla f_{i,\lambda}(\mathbf{w})=(\mathbf{w}-\text{prox}_{\lambda f_{i}}(\mathbf{w}))/\lambda\in-\partial f_{i}(\text{prox}_{\lambda f_{i}}(\mathbf{w}))$, we have $\|\nabla f_{i,\lambda}(\mathbf{w})\|\leq C_{f}$ by Assumption 3.2. Therefore, by Assumption 3.2 again, we can easily show that $\nabla f_{i,\lambda}(g_{i}(\mathbf{w}))\nabla g_{i}(\mathbf{w})$ is $(C_{g}^{2}L_{f_{\lambda}}+C_{f}L_{g})$-Lipschitz continuous (see Lemma 4.2 in [47]), so is $\nabla F_{\lambda}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,\lambda}(g_{i}(\mathbf{w}))\nabla g_{i}(\mathbf{w})$. ∎
@@
$=\min_{\mathbf{y}\in\Omega^{n}}\max_{\mathbf{z}_{i}\in\mathcal{Z}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{y}_{i}^{\top}\mathbf{z}_{i}-f_{i}^{*}(\mathbf{z}_{i})+\frac{1}{2\lambda}\left\|\mathbf{y}_{i}-g_{i}(\mathbf{w})\right\|^{2}$
$=\max_{\mathbf{z}_{i}\in\mathcal{Z}}\min_{\mathbf{y}\in\Omega^{n}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{y}_{i}^{\top}\mathbf{z}_{i}-f_{i}^{*}(\mathbf{z}_{i})+\frac{1}{2\lambda}\left\|\mathbf{y}_{i}-g_{i}(\mathbf{w})\right\|^{2}$
@@
$F_{\lambda}(\mathbf{w})=\max_{\mathbf{z}_{i}\in\mathcal{Z}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{z}_{i}^{\top}g_{i}(\mathbf{w})-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}$
@@
Compare this with the above equality we have
@@
$f_{i,\lambda}(g)=\max_{\mathbf{z}_{i}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}\text{ and }f_{i}(g)=\max_{\mathbf{z}_{i}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})$
@@
$f_{i}(g)=\max_{\mathbf{z}_{i}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})\leq\max_{\mathbf{z}_{i}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}+\max_{\mathbf{z}_{i}}\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}\leq f_{i,\lambda}(g)+\frac{\lambda}{2}C_{f}^{2}$
@@ Appendix C Convergence Analysis of ALEXR2
$F_{\lambda,\nu}(\mathbf{w})$ is $L_{F_{\lambda,\nu}}$-smooth with $L_{F_{\lambda,\nu}}=\frac{2-\nu\rho_{F_{\lambda}}}{\nu-\nu^{2}\rho_{F_{\lambda}}}$ and $\rho_{F_{\lambda}}=\sqrt{d_{1}}C_{f}\rho_{g}$.
@@ C.1 Proof of Technical Lemma
For this lemma we only highlight the difference to the proof of lemma 16 in ALEXR [5]. We define $\Delta_{k}\coloneqq\frac{1}{S}\sum_{i\in\mathcal{S}_{k}}[\nabla g_{i}(\mathbf{w}_{k};\tilde{\mathcal{B}}_{k}^{(i)})]^{\top}y_{k+1}^{(i)}-\frac{1}{n}\sum_{i=1}^{n}[\nabla g_{i}(\mathbf{w}_{k})]^{\top}\bar{y}_{k+1}^{(i)}$.
@@
[…]
@@ C.2 Proof of Inner Loop
where $\tilde{O}(\cdot)$ hides the polylog$(C_{f}C_{g}\rho_{g}/\epsilon^{\prime})$ factor and the green term is the dominant term so $K_{t}=\tilde{O}(\frac{n\sigma_{0}^{2}}{B_{2}B_{1}\epsilon\epsilon^{\prime}})$ ∎
@@ C.3 Proof of Outer loop
$\quad\quad+(1+\frac{2}{\beta})\beta^{2}\left\|\frac{1}{\nu}(\mathbf{w}_{t}-\hat{\mathbf{z}}_{t})-\nabla F_{\lambda,\nu}(\mathbf{w}_{t})\right\|^{2}$
@@
When $g_{i}$ is $L_{g}$-smooth, $F_{\lambda}(\cdot)$ is $(C_{f}L_{g}+C_{g}L_{f})$-smooth. By Theorem 4.10,
@@
[…]
@@
after $K=\tilde{O}\left(\frac{nL_{f}}{B_{1}B_{2}\epsilon^{4}}+\frac{L_{f}}{\epsilon^{4}}\right)$ iterations. Since $C_{f}L_{g}+C_{g}L_{f}=O(\frac{1}{\epsilon})$, we have
$\mathbb{E}\|\nabla F_{\lambda}(\hat{\mathbf{w}}_{\tau})\|\leq\mathbb{E}\|\nabla F_{\lambda}(\text{prox}_{\nu F_{\lambda}}(\mathbf{w}_{\tau}))\|+(C_{f}L_{g}+C_{g}L_{f})\mathbb{E}\left\|\text{prox}_{\nu F_{\lambda}}(\mathbf{w}_{\tau})-\hat{\mathbf{w}}_{\tau}\right\|\leq O(\epsilon),$
[…]
Appendix D Convergence Analysis of ALEXR2 applying to Constraint Optimization Problems
@@
$\nabla f_{\lambda}(\cdot)=\frac{1}{\lambda m}\min\{[\cdot]_{+},\lambda\rho\}.$
Suppose $\mathbf{w}$ is a nearly $\epsilon$-stationary point of (9), which means there exists $\hat{\mathbf{w}}$ such that $\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\text{dist}(0,\partial\Phi_{\lambda}(\hat{\mathbf{w}}))\leq\epsilon$. This means there exist $\mathbf{h}_{i}(\hat{\mathbf{w}})\in\partial g_{i}(\hat{\mathbf{w}})$ and
@@
Suppose $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})>\epsilon$. Then there exists $k$ such that $[g_{k}(\hat{\mathbf{w}})]_{+}>\epsilon$. Recall that $\lambda=\frac{\epsilon}{\rho}$. We have
@@
$\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|\geq\left\|\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|-\left\|\nabla{g}_{0}(\hat{\mathbf{w}})\right\|\geq\sigma_{min}(\mathbf{J}(\hat{\mathbf{w}}))\left\|\bm{\nu}\right\|-\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})\right\|\geq\frac{\delta\rho}{m}-C_{g}>1,$
@@
Finally, when $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq\epsilon$, we have
$\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|=\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\frac{\min\{[g_{i}(\hat{\mathbf{w}})]_{+},\lambda\rho\}}{\lambda m}\leq\frac{\rho}{m}\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\leq O(\epsilon).$ (34)
@@
Suppose Assumptions 3.2(A3) and 5.1 hold. If $\rho>\frac{m(C_{g}+1)}{\delta}$ and $\lambda=\frac{\epsilon}{\rho}$, a nearly $\epsilon$-stationary solution $\mathbf{w}$ to (9) in expectation is also a nearly $O(\epsilon)$-KKT solution to the original problem (7) in the sense that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\mathbb{E}\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\mathbb{E}\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$ and it holds with probability $1-O(\epsilon)$ that $\max\limits_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon)$.
@@
$\nabla f_{\lambda}(\cdot)=\frac{1}{\lambda m}=\min\{[\cdot]_{+},\lambda\rho\}.$
Suppose $\mathbf{w}$ is a nearly $\epsilon$-stationary point of (9) in expecation, which means there exists $\hat{\mathbf{w}}$ such that $\mathbb{E}\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\mathbb{E}\text{dist}(0,\partial\Phi_{\lambda}(\hat{\mathbf{w}}))\leq\epsilon$. This means there exist $\mathbf{h}_{i}(\hat{\mathbf{w}})\in\partial g_{i}(\hat{\mathbf{w}})$ and
@@
$\mathbb{E}\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|\leq\epsilon$ (35)
@@
Suppose $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})>\epsilon$. Then there exists $k$ such that $[g_{k}(\hat{\mathbf{w}})]_{+}>\epsilon$. Recall that $\lambda=\frac{\epsilon}{\rho}$. We have
@@
ϵitalic-$\epsilon\geq$ $\mathbb{E}\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|$
@@
Finally, when $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq\epsilon$, we have
$\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|=\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\frac{\min\{[g_{i}(\hat{\mathbf{w}})]_{+},\lambda\rho\}}{\lambda m}\leq\frac{\rho}{m}\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\leq O(\epsilon).$
It then follows from (36) that
$\text{Prob}\left(\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|\geq O(\epsilon)\right)\leq O(\epsilon).$ (37)
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+g_{0}(\mathbf{z}_{k+1})-g_{0}(\mathbf{w})-\left\langle G_{k}^{0},\,\mathbf{z}_{k+1}-\mathbf{z}\right\rangle},$
@@
$g_{0}(\mathbf{w})-g_{0}(\mathbf{z}_{k})\geq\left\langle g_{0}^{\prime}(\mathbf{z}_{k}),\,\mathbf{w}-\mathbf{z}_{k}\right\rangle-\frac{\rho_{g}}{2}\|\mathbf{z}-\mathbf{z}_{k}\|_{2}^{2}.$

### v2

@@ 7 Conclusions and Discussion
Acknowledgments and Disclosure of Funding
We thank Quanqi Hu for the assistance in developing the codebase for group DRO. X. Chen, B. Wang, M. Yang and T. Yang were partially supported by the National Science Foundation Career Award 2246753, the National Science Foundation Award 2246757 and 2306572. Q. Lin was partially supported by National Science Foundation Award 2147253.
@@ Appendix A Proofs of Technical Lemmas
Under Assumption 3.2(A1), $F(\cdot)$ in (1) is $\rho_{F}$-weakly convex with $\rho_{F}=\sqrt{d_{1}}L_{g}C_{f}+\rho_{f}C_{g}^{2}$. If $g_{i}$ is $L_{g}$ smooth, then $F_{\lambda}(\mathbf{w})$ is $L_{F}$-smooth with $L_{F}=C_{g}^{2}\max\{\frac{1}{\lambda},\frac{\rho_{f}}{1-\lambda\rho_{f}}\}+C_{f}L_{g}$.
@@
Consider any $\mathbf{w}$ and $\mathbf{w}^{\prime}$ in $\mathbb{R}^{d}$, any $i\in\{1,\dots,n\}$ and any $\mathbf{v}_{i}\in\partial f_{i}(\mathbf{t}_{i})$ at $\mathbf{t}_{i}=g_{i}(\mathbf{w})$. By Assumption 3.2(A1), $\mathbf{v}_{i}^{\top}g_{i}(\mathbf{w})$ is $\|\mathbf{v}_{i}\|_{1}L_{g}$-smooth in $\mathbf{w}$ and $f_{i}$ is $\rho_{f}$-weakly convex, we have
@@
$\geq f_{i}(g_{i}(\mathbf{w}))+\mathbf{v}_{i}^{\top}\nabla g_{i}(\mathbf{w}^{\prime})(\mathbf{w}^{\prime}-\mathbf{w})-\frac{\|\mathbf{v}_{i}\|_{1}L_{g}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2}-\frac{\rho_{f}C_{g}^{2}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2}$
$\geq f_{i}(g_{i}(\mathbf{w}))+\mathbf{v}_{i}^{\top}\nabla g_{i}(\mathbf{w}^{\prime})(\mathbf{w}^{\prime}-\mathbf{w})-\frac{\sqrt{d_{1}}C_{f}L_{g}+\rho_{f}C_{g}^{2}}{2}\left\|\mathbf{w}^{\prime}-\mathbf{w}\right\|^{2},$
where the second and the last inequalities hold because of lipschitzness of $f_{i}$ and $g_{i}$ from Assumption 3.2, the third because $\mathbf{v}_{i}^{\top}g_{i}(\mathbf{w})$ is $\|\mathbf{v}_{i}\|_{1}L_{g}$-smooth in $\mathbf{w}$.
By the $\rho_{f}$-weak convexity of $f_{i}$ and lemma 13 in [33], we know that $f_{i,\lambda}$ is $L_{f_{\lambda}}:=\max\{\frac{1}{\lambda},\frac{\rho_{f}}{1-\lambda\rho_{f}}\}$-smooth. Since $\nabla f_{i,\lambda}(\mathbf{w})=(\mathbf{w}-\text{prox}_{\lambda f_{i}}(\mathbf{w}))/\lambda\in\partial f_{i}(\text{prox}_{\lambda f_{i}}(\mathbf{w}))$, we have $\|\nabla f_{i,\lambda}(\mathbf{w})\|\leq C_{f}$ by Assumption 3.2. Therefore, by Assumption 3.2 again, we can easily show that $\nabla f_{i,\lambda}(g_{i}(\mathbf{w}))\nabla g_{i}(\mathbf{w})$ is $(C_{g}^{2}L_{f_{\lambda}}+C_{f}L_{g})$-Lipschitz continuous (see Lemma 4.2 in [44]), so is $\nabla F_{\lambda}(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i,\lambda}(g_{i}(\mathbf{w}))\nabla g_{i}(\mathbf{w})$. ∎
@@
$=\min_{\mathbf{y}\in\Omega^{n}}\max_{\mathbf{z}\in\mathcal{Z}^{n}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{y}_{i}^{\top}\mathbf{z}_{i}-f_{i}^{*}(\mathbf{z}_{i})+\frac{1}{2\lambda}\left\|\mathbf{y}_{i}-g_{i}(\mathbf{w})\right\|^{2}$
$=\max_{\mathbf{z}\in\mathcal{Z}^{n}}\min_{\mathbf{y}\in\Omega^{n}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{y}_{i}^{\top}\mathbf{z}_{i}-f_{i}^{*}(\mathbf{z}_{i})+\frac{1}{2\lambda}\left\|\mathbf{y}_{i}-g_{i}(\mathbf{w})\right\|^{2}$
@@
$F_{\lambda}(\mathbf{w})=\max_{\mathbf{z}\in\mathcal{Z}^{n}}\frac{1}{n}\sum_{i=1}^{n}\mathbf{z}_{i}^{\top}g_{i}(\mathbf{w})-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}$ (12)
@@
Compare this with 12 we have
@@
$f_{i,\lambda}(g)=\max_{\mathbf{z}_{i}\in\mathcal{Z}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}\text{ and }f_{i}(g)=\max_{\mathbf{z}_{i}\in\mathcal{Z}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})$
@@
$f_{i}(g)=\max_{\mathbf{z}_{i}\in\mathcal{Z}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})\leq\max_{\mathbf{z}_{i}\in\mathcal{Z}}\mathbf{z}_{i}^{\top}g-f_{i}^{*}(\mathbf{z}_{i})-\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}+\max_{\mathbf{z}_{i}\in\mathcal{Z}}\frac{\lambda}{2}\left\|\mathbf{z}_{i}\right\|^{2}\leq f_{i,\lambda}(g)+\frac{\lambda}{2}C_{f}^{2}$
@@ Appendix C Convergence Analysis of ALEXR2
$F_{\lambda,\nu}(\mathbf{w})$ is $L_{F_{\lambda,\nu}}$-smooth with $L_{F_{\lambda,\nu}}=\max\{\frac{1}{\nu},\frac{\rho_{F_{\lambda}}}{1-\nu\rho_{F_{\lambda}}}\}$ and $\rho_{F_{\lambda}}=\sqrt{d_{1}}C_{f}\rho_{g}$.
@@ C.1 Proof of Technical Lemma
For this lemma we only highlight the difference to the proof of lemma 7 in ALEXR [5]. We define $\Delta_{k}\coloneqq\frac{1}{B_{1}}\sum_{i\in\mathcal{B}_{1}^{k}}[\nabla g_{i}(\mathbf{w}_{k};\tilde{\mathcal{B}}_{i,2}^{k})]^{\top}y_{k+1}^{(i)}-\frac{1}{n}\sum_{i=1}^{n}[\nabla g_{i}(\mathbf{w}_{k})]^{\top}\bar{y}_{k+1}^{(i)}$.
@@
$\quad\quad+\frac{L_{g}}{2}\left\|\mathbf{w}_{k}-\mathbf{w}^{*}\right\|^{2}*\frac{1}{n}\sum_{i=1}^{n}\left\langle\bar{y}_{k+1}^{(i)},\,\mathbf{1}\right\rangle+\left\langle\frac{1}{n}\sum_{i=1}^{n}[\nabla g_{i}(\mathbf{w}_{k})]^{\top}\bar{y}_{k+1}^{(i)}+\Delta_{k},\,\mathbf{w}_{*}-\mathbf{w}_{k+1}\right\rangle$
@@ C.2 Proof of Inner Loop
[…]
@@ C.3 Proof of Outer loop
$\qquad\qquad+(1+\frac{2}{\beta})\beta^{2}\left\|\frac{1}{\nu}(\mathbf{w}_{t}-\hat{\mathbf{z}}_{t})-\nabla F_{\lambda,\nu}(\mathbf{w}_{t})\right\|^{2}$
@@
When $g_{i}$ is $L_{g}$-smooth, $F_{\lambda}(\cdot)$ is $L_{F}$-smooth where $L_{F}=O(\frac{1}{\lambda})=O(\frac{1}{\epsilon})$ is defined in lemma A.1. By Theorem 4.10,
@@
According to Theorem C.1, for $\epsilon^{\prime}=O(\epsilon^{4})$, there exists $\theta\in(0,1)$ with $1-\theta=O(\epsilon^{4})$ such that, by setting $\eta=\frac{1-\theta}{\theta}L_{f}$ and $\gamma=\frac{(1-\theta)n}{B_{1}}$ in inner loop of Algorithm 2(i.e. ALEXR), $\hat{\mathbf{w}}_{\tau}=$ALEXR($\mathbf{w}_{\tau},K)$ satisfies $\frac{\mu}{2}\mathbb{E}\left\|\hat{\mathbf{w}}_{\tau}-\text{prox}_{\nu F_{\lambda}}(\mathbf{w}_{\tau})\right\|^{2}\leq\epsilon^{4}$ and thus
@@
after $K=\tilde{O}\left(\frac{nL_{f}}{B_{1}B_{2}\epsilon^{4}}+\frac{L_{f}}{\epsilon^{4}}\right)$ iterations. Then we have
$\mathbb{E}\|\nabla F_{\lambda}(\hat{\mathbf{w}}_{\tau})\|\leq\mathbb{E}\|\nabla F_{\lambda}(\text{prox}_{\nu F_{\lambda}}(\mathbf{w}_{\tau}))\|+L_{F}\mathbb{E}\left\|\text{prox}_{\nu F_{\lambda}}(\mathbf{w}_{\tau})-\hat{\mathbf{w}}_{\tau}\right\|\leq O(\epsilon),$
which means $\hat{\mathbf{w}}_{\tau}$ is an $\epsilon$-stationary solution of (3). By Proposition 4.4, $\hat{\mathbf{w}}_{\tau}$ is a nearly $\epsilon$-stationary solution of (1). By Theorem 4.10, $\mathbf{w}_{\tau}$ is found within complexity $O(\epsilon^{-5})$ and $\hat{\mathbf{w}}_{\tau}=$ALEXR($\mathbf{w}_{\tau},K)$ has complexity $K=\tilde{O}\left(\frac{nL_{f}}{B_{1}B_{2}\epsilon^{4}}+\frac{L_{f}}{\epsilon^{4}}\right)$. The total complexity for computing $\hat{\mathbf{w}}_{\tau}$ is still $\tilde{O}\left(\frac{n}{B_{1}B_{2}\epsilon^{5}}+\frac{1}{\epsilon^{5}}\right)$. ∎
Appendix D Convergence Analysis of SONEX and ALEXR2 applying to Constraint Optimization Problems
@@
$\nabla f_{\lambda}(\cdot)=\frac{1}{\lambda}\min\{[\cdot]_{+},\lambda\rho\}.$
[…]
@@
Suppose $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})>\epsilon$. Then there exists $k$ such that $g_{k}(\hat{\mathbf{w}})]_{+}>\epsilon$. Recall that $\lambda=\frac{\epsilon}{\rho}$. We have
@@
$\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|\geq\left\|\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|-\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})\right\|\geq\sigma_{min}(\mathbf{J}(\hat{\mathbf{w}}))\left\|\bm{\nu}\right\|-\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})\right\|\geq\frac{\delta\rho}{m}-C_{g}>1,$
@@
Finally, when $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq\epsilon$, we have for $\forall i=1,2,\cdots,m$
$|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|=\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\frac{\min\{[g_{i}(\hat{\mathbf{w}})]_{+},\lambda\rho\}}{\lambda m}\leq\frac{\rho}{m}\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\leq O(\epsilon).$ (35)
@@
Suppose Assumptions 3.2(A3) and 5.1 hold. If $\rho>\frac{m(C_{g}+1)}{\delta}$ and $\lambda=\frac{\epsilon}{\rho}$, a nearly $\epsilon$-stationary solution $\mathbf{w}$ to (9) in expectation is also a nearly $O(\epsilon)$-KKT solution to the original problem (7) in the sense that there exist $\hat{\mathbf{w}}$ and $\nu_{i}\geq 0$ for $i=1,\dots,m$ such that $\mathbb{E}\left\|\mathbf{w}-\hat{\mathbf{w}}\right\|\leq\epsilon$ and $\mathbb{E}\text{dist}(0,\partial g_{0}(\hat{\mathbf{w}})+\sum_{i=1}^{m}\partial g_{i}(\hat{\mathbf{w}})\nu_{i}))\leq O(\epsilon)$ and it holds with probability $1-O(\epsilon)$ that $\max\limits_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq O(\epsilon)$ and $|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq O(\epsilon),\forall i=1,2,\cdots,m$.
@@
$\nabla f_{\lambda}(\cdot)=\frac{1}{\lambda}=\min\{[\cdot]_{+},\lambda\rho\}.$
[…]
@@
$\mathbb{E}\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|\leq\epsilon$
@@
Suppose $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})>\epsilon$. Then there exists $k$ such that $g_{k}(\hat{\mathbf{w}})]_{+}>\epsilon$. Recall that $\lambda=\frac{\epsilon}{\rho}$. We have
@@
$\epsilon\geq$ $\mathbb{E}\left\|\mathbf{h}_{0}(\hat{\mathbf{w}})+\mathbf{J}(\hat{\mathbf{w}})\bm{\nu}\right\|$
@@
Finally, when $\max_{i=1,\dots,m}g_{i}(\hat{\mathbf{w}})\leq\epsilon$, we have for $\forall i=1,2,\cdots,m$
$|g_{i}(\hat{\mathbf{w}})\nu_{i}|\leq\sum_{i=1}^{m}|g_{i}(\hat{\mathbf{w}})\nu_{i}|=\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\frac{\min\{[g_{i}(\hat{\mathbf{w}})]_{+},\lambda\rho\}}{\lambda m}\leq\frac{\rho}{m}\sum_{i=1}^{m}[g_{i}(\hat{\mathbf{w}})]_{+}\leq O(\epsilon).$
It then follows from (36) that for $\forall i=1,2,\cdots,m$
$\text{Prob}\left(|g_{i}(\hat{\mathbf{w}})\nu_{i}|\geq O(\epsilon)\right)\leq O(\epsilon).$
@@
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+g_{0}(\mathbf{z}_{k+1})-g_{0}(\mathbf{z})-\left\langle G_{k}^{0},\,\mathbf{z}_{k+1}-\mathbf{z}\right\rangle},$
@@
$g_{0}(\mathbf{z})-g_{0}(\mathbf{z}_{k})\geq\left\langle g_{0}^{\prime}(\mathbf{z}_{k}),\,\mathbf{z}-\mathbf{z}_{k}\right\rangle-\frac{\rho_{g}}{2}\|\mathbf{z}-\mathbf{z}_{k}\|_{2}^{2}.$